# Exact Gauss-Newton Optimization for Training Deep Neural Networks

### v1

@@ 3 Algorithm
We are now ready to present the EGN algorithm. First, we will discuss how one can efficiently solve the linear system. Then, we will discuss further enhancements to the basic algorithm.
Next, we address issue (c), i.e., the problem of finding the solution of the symmetric linear system $\mathbf{H}_{t}\mathbf{d}_{t}=-\mathbf{g}_{t}$. Substituting the exact Hessian with the regularized Gauss-Newton Hessian yields
@@
To that end, we utilize the Duncan-Guttman identity, originally proven in [23, 26].
Theorem 3.1
@@
By observing that Equation (3.1) defining the direction $\mathbf{d}^{\mathrm{LM}}$ has the form of the left-hand side of the identity (3.2) we state the following (see proof in Appendix B.4).
@@
The key property of Algorithm 1 is that the system (2.5) is solved exactly in contrast to approximate (or inexact) solutions that underpin algorithms such as HFO [46, 37], Newton-SGI [6], LiSSA [2], SGN [24] and iGN [70]. The benefits of having the exact solution are analyzed in [6] with exact Newton methods enjoying faster convergence rates than inexact ones. Our experiments (Section 5) also demonstrate that the exact Gauss-Newton solver (EGN) consistently outperforms the inexact version (SGN) across the majority of the problems. The complexity of Algorithm 1 is dominated by the matrix multiplication $\mathbf{J}\mathbf{J}^{\top}$ that costs $\mathcal{O}\left(b^{2}c^{2}d\right)$ as well as solving the linear system of size $bc\times bc$ with complexity $\mathcal{O}\left(b^{3}c^{3}\right)$. Assuming $d>bc$, the overall complexity of Algorithm 1 is $\mathcal{O}\left(b^{2}c^{2}d\right)$.
An alternative approach to solve Equation (3.1) exactly is the SWM formula employed in [60, 13, 31, 4]. Solving (3.1) with the SWM formula yields $\mathbf{d}_{t}=-\frac{1}{\lambda}\left(\mathbf{I}_{d}-\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\left(\lambda\mathbf{I}_{bc}+\frac{1}{b}\mathbf{J}_{t}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\right)^{-1}\mathbf{J}_{t}\right)\mathbf{g}_{t}$. As with EGN, the dominating term is $\mathbf{J}\mathbf{J}^{\top}$. However, even with efficient ordering of operations, there are at least two $\mathcal{O}\left(b^{2}c^{2}d\right)$ matrix multiplications while EGN requires just one. Another important consideration is the non-regularized case, i.e., when $\lambda=0$. Assuming $\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}$ is invertible, EGN produces a valid solution while SWM is not defined in that case. That makes Algorithm 1 a preferred choice for finding the direction (B.6) exactly.
@@
For completeness, we note here that another option to solve (3.1) exactly is through a QR factorization of $\mathbf{J}^{\top}_{t}$ (see Appendix C). The complexity of such an approach is also $\mathcal{O}\left(b^{2}c^{2}d\right)$, dominated by the economy size QR decomposition of $\mathbf{J}^{\top}$. However, performing a QR decomposition is significantly more expensive than performing matrix-matrix multiplications and, in practice, EGN is preferred.
@@
The CG method is an essential part of, e.g., HFO [46, 37], SGN [24], Newton-CG [6], LM [59], and Distributed Newton’s method with optimal shrinkage [82]. The complexity of CG approximately solving Equation (3.1) is $\mathcal{O}(ld^{2})$ where $l$ is the number of CG iterations [53]. A typical number of CG iterations ranges from $3$ in [37] to $50$ in [15]. Note that, unless the number of classes $c$ is high (which is one of the limitations of EGN addressed in Section 5.3), we have $ld^{2}>b^{2}c^{2}d$, such that EGN solves the system both exactly and faster than CG.
@@
Besides estimating the Hessian-adjusted direction, most state-of-the-art SOMs employ strategies to reduce variance of gradient and Hessian estimates, safeguard against exploding gradients, and dynamically adjust hyper-parameters to training steps. In Appendix C we discuss several enhancements to EGN, including momentum acceleration [36, 37, 78] to mitigate issue (d), i.e., the noise in Hessian estimates; line search [17, 72, 59] to ensure steps are sufficiently short to decrease the loss; and adaptive regularization [37, 59, 60] to obtain faster convergence and simplify hyper-parameter tuning. The pseudocode for EGN is outlined in Algorithm 5 in Appendix C.
@@ 4 Convergence Analysis
In this section, we analyze the convergence of EGN (Algorithm 1) in the general non-convex setting. In EGN, we consider the sequence of iterates $\{\mathbf{w}_{t}\}_{t\geq 1}$ where each $\mathbf{w}_{t}$ is computed via (2.4) with $\mathbf{d}_{t}\equiv\mathbf{d}_{t}^{\mathrm{LM}}$. We aim to minimize the function ${\mathcal{L}}_{N}(\mathbf{w})$ using, for each realization $\xi\sim P_{\xi}$, the Hessian estimator $\mathbf{H}(\mathbf{w},\xi)$ and the gradient estimator $\mathbf{g}(\mathbf{w},\xi)$. Then, at each iteration $t$, the mini-batch estimates $\mathbf{g}_{t}$ and $\mathbf{H}_{t}$ of the gradient and Hessian are
$\mathbf{g}_{t}=\frac{1}{b}\sum_{\xi_{i}\in{\mathcal{B}}_{t}}\mathbf{g}(\mathbf{w},\xi_{i})\coloneqq\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{r}_{t},\quad\mathbf{H}_{t}=\frac{1}{b}\sum_{\xi_{i}\in{\mathcal{B}}_{t}}\mathbf{H}(\mathbf{w},\xi_{i})\coloneqq\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}.$ (4.1)
@@
The function ${\mathcal{L}}_{N}$ is lower-bounded on its domain, i.e., $-\infty<{\mathcal{L}}_{N}^{*}\coloneqq\inf\limits_{\mathbf{w}\in{\mathbb{R}}^{d}}{\mathcal{L}}_{N}(\mathbf{w})$. In addition, ${\mathcal{L}}_{N}$ is twice differentiable with Lipschitz continuous first-order derivatives, i.e., $\exists L_{1}\in\mathbb{R}$ such that
@@
At any iteration $t$, $\underaccent{\bar}{\kappa}\mathbf{I}\leq\mathbf{Q}_{t}\leq\bar{\kappa}\mathbf{I}$ with $\bar{\kappa}\geq\underaccent{\bar}{\kappa}\geq 0$. Moreover, $\exists\bar{\sigma},\underaccent{\bar}{\sigma}$ satisfying $0<\underaccent{\bar}{\sigma}\leq\bar{\sigma}$ such that $\underaccent{\bar}{\sigma}\leq\left\lVert\mathbf{J}_{t}\right\rVert\leq\bar{\sigma}$ for all $t\geq 0$.
@@
In Lemma 4.1, we prove a descent lemma for EGN under the given conditions (see Appendix B.5 for the proof).
@@
Let $\{\mathbf{w}_{t}\}$ be the sequence of iterates generated by (2.4) with $\mathbf{d}_{t}\equiv\mathbf{d}_{t}^{\mathrm{LM}}$ and let Assumptions 4.14.3 hold. Assume further that $\alpha_{t}\leq(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}/(b^{2}\lambda_{t}L_{1})$ holds. Then,
@@
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\frac{\alpha_{t}}{10\lambda_{t}}\left\lVert\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\right\rVert^{2}+\frac{\sigma_{g}^{2}\alpha_{t}}{2b\lambda_{t}}.$ (4.5)
@@
Assumption 4.4 is necessary to prove the result in Lemma 4.1 due to the correlation between $\mathbf{B}_{t}$ and $\mathbf{g}_{t}$. If these quantities were not correlated, then one could exploit Assumptions 4.24.3: by taking the expected value and using (4.3), one could then bound $\mathbf{B}_{t}\succeq\mu\mathbf{I}$ to obtain the desired bound with $K=1$. Clearly, that would yield different constants in our next results, but their nature would remain unaltered. Note that this observation could suggest alternative versions of EGN, in which the correlation is removed by construction at the expense of an increased computational burden. A thorough investigation of such schemes is beyond the scope of this paper.
We next show the iteration complexity estimate of EGN to reach an $\epsilon$-stationary point. We provide the proof in Appendix B.6.

### v2

@@ 4 Algorithm
We are now ready to present the EGN algorithm. First, we will discuss how one can efficiently solve the linear system. Then, we will discuss further enhancements to the basic algorithm. Next, we address issue (c), i.e., the problem of finding the solution of the symmetric linear system $\mathbf{H}_{t}\mathbf{d}_{t}=-\mathbf{g}_{t}$.
Substituting the exact Hessian with the regularized Gauss-Newton Hessian yields
@@
To that end, we follow an approach similar to the ones in adeoye2023score ; adeoye2021sc which propose to utilize the Duncan-Guttman identity duncan1944lxxviii ; guttman1946enlargement . We present this in Theorem 4.1 and Lemma 4.2 below for the Levenberg-Marquardt direction. For a general presentation which considers smooth regularization functions, we refer the interested reader to (adeoye2023score, , Section 3).
Theorem 4.1 (duncan1944lxxviii ; guttman1946enlargement )
@@
By observing that Equation (4.1) defining the direction $\mathbf{d}^{\mathrm{LM}}$ has the form of the left-hand side of the identity (4.2) we state the following.
@@
$\left(b\lambda\mathbf{I}_{d}+\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\mathbf{Q}=-\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\left(\mathbf{Q}^{-1}+\mathbf{J}\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\right)^{-1}$ (4.3)
Multiplying both sides by $\mathbf{Q}^{-1}\mathbf{r}$ yields
$\left(b\lambda\mathbf{I}_{d}+\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\mathbf{r}=-\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\left(\mathbf{Q}^{-1}+\mathbf{J}\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}\mathbf{r}$ (4.4)
where the lhs of (4.4) is now equivalent to explicitly solving for $\mathbf{d}^{\mathrm{LM}}$ in Equation (4.1). Simplifying the rhs of (4.4) results in
$\mathbf{d}^{\mathrm{LM}}=-\mathbf{J}^{\top}\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}\mathbf{r}.$ (4.5)
$\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}=\left(\mathbf{Q}\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)\right)^{-1}=\left(b\lambda\mathbf{I}_{bc}+\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}\right)^{-1}.$ (4.6)
So that the direction can be calculated as
$\mathbf{d}^{\mathrm{LM}}=-\mathbf{J}^{\top}\left(\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}+b\lambda\mathbf{I}_{bc}\right)^{-1}\mathbf{r},$ (4.7)
which corresponds to the procedure outlined in Algorithm 1. $\Box$
4.1 Comparison to Existing Methods
@@
The key property of Algorithm 1 is that the system (3.5) is solved exactly in contrast to approximate (or inexact) solutions that underpin algorithms such as HFO martens2010deep ; kiros2013training , Newton-SGI bollapragada2019exact , LiSSA agarwal2017second , SGN gargiani2020promise and iGN tran2020stochastic . The benefits of having the exact solution are analyzed in bollapragada2019exact with exact Newton methods enjoying faster convergence rates than inexact ones. Our experiments (Section 6) also demonstrate that the exact Gauss-Newton solver (EGN) consistently outperforms the inexact version (SGN) across the majority of the problems. The complexity of Algorithm 1 is dominated by the matrix multiplication $\mathbf{J}\mathbf{J}^{\top}$ that costs $\mathcal{O}\left(b^{2}c^{2}d\right)$ as well as solving the linear system of size $bc\times bc$ with complexity $\mathcal{O}\left(b^{3}c^{3}\right)$. Assuming $d>bc$, the overall complexity of Algorithm 1 is $\mathcal{O}\left(b^{2}c^{2}d\right)$.
A common alternative way to solve the Levenberg-Marquardt linear system (4.1) exactly is to apply the SMW identity, as in ren2019efficient ; brust2021nonlinear ; hong2020stochastic ; arbel2024rethinking . The SMW identity states
$\left(\mathbf{A}+\mathbf{U}\mathbf{C}\mathbf{V}\right)^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\mathbf{U}\left(\mathbf{C}^{-1}+\mathbf{V}\mathbf{A}^{-1}\mathbf{U}\right)^{-1}\mathbf{V}\mathbf{A}^{-1}.$ (4.8)
With $\mathbf{A}=\lambda\mathbf{I}_{d}$, $\mathbf{U}=\mathbf{J}^{\top}$, $\mathbf{C}=\frac{1}{b}\mathbf{Q}$, and $\mathbf{V}=\mathbf{J}$, we obtain the SMW-GN matrix inversion ren2019efficient :
$\left(\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}+\lambda\mathbf{I}_{d}\right)^{-1}=\frac{1}{\lambda}\mathbf{I}_{d}-\frac{1}{\lambda^{2}}\mathbf{J}^{\top}\left(b\mathbf{Q}^{-1}+\frac{1}{\lambda}\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{J}.$ (4.9)
As with EGN, the dominant term is $\mathbf{J}\mathbf{J}^{\top}$. However, even with efficient ordering of operations, there are at least two $\mathcal{O}\left(b^{2}c^{2}d\right)$ matrix multiplications while EGN requires just one. Beyond doubling the per-step cost, the additional multiplication can introduce extra rounding error that gradually accumulates higham2002accuracy .
Another important consideration is the non-regularized case. When Levenberg damping is disabled ($\lambda=0$), $\mathbf{A}$ in (4.8) is singular and the SMW inverse is undefined. EGN remains well-posed provided $\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}$ is invertible, making it applicable to pure Gauss-Newton steps.
We support our theoretical findings with the empirical comparison of solving Equation (4.1) using both SMW and EGN methods (Table 2). EGN achieves up to $1.6\times$ speed-up compared to SMW on larger models ($d\!\geq\!10^{5}$), which translates into substantial training-time savings since the solver is invoked at every iteration.
Table 2: Wall-clock time (seconds, mean over 1000 runs) to solve (4.1) with SMW and EGN across different model sizes for $b=32$, $c=10$ on NVIDIA RTX A4000 GPU.
$d$ 1K 10K 100K 1M 2M
SMW 0.0010 0.0013 0.0044 0.0368 0.0747
EGN 0.0008 0.0011 0.0028 0.0246 0.0517
@@
For completeness, we note here that another option to solve (4.1) exactly is through the QR factorization of $\mathbf{J}^{\top}_{t}$ (see pseudocode in Appendix B). The complexity of such an approach is also $\mathcal{O}\left(b^{2}c^{2}d\right)$, dominated by the economy size QR decomposition of $\mathbf{J}^{\top}$. However, performing a QR decomposition is significantly more expensive than performing matrix-matrix multiplications and, in practice, EGN is preferred.
@@
The CG method is an essential part of, e.g., HFO martens2010deep ; kiros2013training , SGN gargiani2020promise , Newton-CG bollapragada2019exact , LM pooladzandi2022improving , and Distributed Newton’s method with optimal shrinkage zhang2023optimal . The complexity of CG approximately solving Equation (4.1) is $\mathcal{O}(ld^{2})$ where $l$ is the number of CG iterations nocedal2006numerical . A typical number of CG iterations ranges from $3$ in kiros2013training to $50$ in chapelle2011improved . Note that, unless the number of classes $c$ is high (which is one of the limitations of EGN addressed in Section 6.3), we have $ld^{2}>b^{2}c^{2}d$, such that EGN solves the system both exactly and faster than CG.
4.2 Additional Improvements
@@
In addition to estimating the Hessian-adjusted direction, most state-of-the-art SOMs employ strategies to reduce variance of gradient and Hessian estimates, safeguard against exploding gradients, and dynamically adjust hyper-parameters to training steps. Next, we explore several enhancements to EGN, including momentum acceleration kingma2014adam ; kiros2013training ; yao2021adahessian to mitigate issue (d)–that is, the noise in Hessian estimates; line search curtis2020adaptive ; vaswani2019painless ; pooladzandi2022improving to ensure steps are sufficiently short to decrease the loss; and adaptive regularization kiros2013training ; pooladzandi2022improving ; ren2019efficient to achieve faster convergence and simplify hyper-parameter tuning.
Momentum
A significant challenge in second-order methods is the noise introduced in Hessian estimates due to stochastic sampling. Consider the update direction $\mathbf{d}_{t}=-\mathbf{H}^{-1}_{t}\mathbf{g}_{t}$, where both the approximate Hessian $\mathbf{H}_{t}$ and the gradient $\mathbf{g}_{t}$ are stochastic estimates of the true derivatives of the empirical risk (3.2). Conditioning the gradient with a noisy Hessian inverse can lead to inaccurate descent directions, impeding convergence. To mitigate this issue, temporal averaging techniques (or momentum) are employed to stabilize updates and accelerate convergence by combining information from previous iterations with current estimates. In ML applications, common momentum variants include simple accumulation duchi2011adaptive , exponential moving average (EMA) hinton2012neural , bias-corrected EMA kingma2014adam ; yao2021adahessian and momentum with an extrapolation step nesterov1983method .
For diagonal scaling methods, including first-order accelerated algorithms kingma2014adam ; duchi2011adaptive ; hinton2012neural and SOMs yao2021adahessian ; liu2023sophia , the accumulated estimates of both first and second moments are kept separately resulting in $\mathcal{O}\left(d\right)$ space complexity. For Gauss-Newton methods we could alternatively reduce the variance of the Jacobian $\mathbf{J}$, e.g., with SVRG johnson2013accelerating , SAGA defazio2014saga or SARAH nguyen2017sarah , which results in a space complexity of $\mathcal{O}\left(bcd\right)$. Another option is to apply momentum to the descent direction $\mathbf{d}_{t}$, explored in vaswani2019painless ; kiros2013training , which requires storing a vector of size $d$. Since we never explicitly materialize either the Hessian or the gradient (Algorithm 1), we follow the latter approach with bias-corrected EMA by default.
Algorithm 2 EGN
2: Initialize momentum: $\mathbf{m}_{0}=0$
3:for $t$ in $1..T$ do
4:  Sample a mini-batch $\mathcal{B}_{t}$ from $\mathcal{D}$
5:  Estimate $\mathbf{r}_{t}$ and $\mathbf{J}_{t}$ (e.g., via backpropagation)
6:  Find direction $\mathbf{d}_{t}$ via Algorithm 1
9:  Line search for $\alpha_{t}$ via Algorithm 4
11:  Update $\lambda_{t+1}$ via Algorithm 5
12:end for
13:Return $\mathbf{w}_{t}$
Line Search
Line search is a widely used technique in deterministic optimization nocedal2006numerical that iteratively adjusts the learning rate to satisfy some minimum criteria (e.g., Wolfe’s conditions), ensuring adequate decrease in the loss function. Allowing $\alpha$ to automatically adapt to each training step can significantly reduce the need for manual tuning, which is often time-consuming and computationally expensive. However, extending line search methods to the stochastic setting poses challenges due to the inherent noise in gradient estimates, making it difficult to guarantee the same theoretical properties as in the deterministic case curtis2020adaptive . Despite these challenges, line search has been successfully applied in practice to both stochastic FOMs vaswani2019painless ; paquette2020stochastic and SOMs pooladzandi2022improving ; wills2021stochastic . In EGN we adopt the strategy proposed by vaswani2019painless , which incorporates a reset mechanism at the beginning of each search to minimize the computational overhead of evaluating the loss function (Algorithm 4 in the Appendix).
Adaptive regularization
Adaptive regularization techniques for stochastic SOMs are explored in martens2010deep ; kiros2013training ; ren2019efficient ; pooladzandi2022improving modifying the original Levenberg-Marquardt rule to the stochastic setting. The central idea is to track $\rho$, defined as the ratio between the decrease in the actual loss function (3.3) and the decrease in the quadratic model $\mathcal{M}\left(\Delta\mathbf{w}\right)=\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t})\right)+\mathbf{g}_{t}^{\top}\Delta\mathbf{w}+\frac{1}{2b}\Delta\mathbf{w}^{\top}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}\Delta\mathbf{w}$, where $\Delta\mathbf{w}=\mathbf{w}-\mathbf{w}_{t}$. This yields
$\rho:=\frac{\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t+1})\right)-\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t})\right)}{\mathbf{g}_{t}^{\top}\Delta\mathbf{w}+\frac{1}{2b}\Delta\mathbf{w}^{\top}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}\Delta\mathbf{w}},$ (4.10)
which measures the accuracy of the quadratic model. In case $\rho$ is small or negative, $\mathcal{M}\left(\Delta\mathbf{w}\right)$ provides inaccurate approximation and the value $\lambda$ is increased. Conversely, if $\rho$ is large, $\lambda$ is decreased to give more weight to $\mathcal{M}\left(\Delta\mathbf{w}\right)$ (see Algorithm 5 in the Appendix). As empirically found by kiros2013training , compared to the deterministic LM the increase/decrease coefficients need to be less aggressive to reduce the oscillations of $\lambda_{t}$.
@@ 5 Convergence Analysis
In this section, we analyze the convergence of EGN in the general non-convex setting.
In EGN, we consider the sequence of iterates $\{\mathbf{w}_{t}\}_{t\geq 1}$ where each $\mathbf{w}_{t}$ is computed via (3.4) with $\mathbf{d}_{t}\equiv\mathbf{d}_{t}^{\mathrm{LM}}$. We aim to minimize the function ${\mathcal{L}}_{N}(\mathbf{w})$ using, for each realization $\xi\sim P_{\xi}$, the Hessian estimator $\mathbf{H}(\mathbf{w},\xi)$ and the gradient estimator $\mathbf{g}(\mathbf{w},\xi)$. Then, at each iteration $t$, the mini-batch estimates $\mathbf{g}_{t}$ and $\mathbf{H}_{t}$ of the gradient and Hessian are
$\mathbf{g}_{t}=\frac{1}{b}\sum_{\xi_{i}\in{\mathcal{B}}_{t}}\mathbf{g}(\mathbf{w},\xi_{i})\coloneqq\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{r}_{t},$ (5.1)
$\mathbf{H}_{t}=\frac{1}{b}\sum_{\xi_{i}\in{\mathcal{B}}_{t}}\mathbf{H}(\mathbf{w},\xi_{i})\coloneqq\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}.$ (5.2)
@@
The function ${\mathcal{L}}_{N}$ is lower-bounded on its domain, i.e., $\infty<{\mathcal{L}}_{N}^{*}\coloneqq\inf\limits_{\mathbf{w}\in{\mathbb{R}}^{d}}{\mathcal{L}}_{N}(\mathbf{w})$. In addition, ${\mathcal{L}}_{N}$ is twice differentiable with Lipschitz continuous first-order derivatives, i.e., $\exists L_{1}\in\mathbb{R}$ such that
@@
At any iteration $t$, $\underaccent{\bar}{\kappa}\mathbf{I}\leq\mathbf{Q}_{t}\leq\bar{\kappa}\mathbf{I}$ with $\bar{\kappa}\geq\underaccent{\bar}{\kappa}\geq 0$. Additionally, $\exists\bar{\sigma},\underaccent{\bar}{\sigma}$ satisfying $0<\underaccent{\bar}{\sigma}\leq\bar{\sigma}$ such that $\underaccent{\bar}{\sigma}\leq\left\lVert\mathbf{J}_{t}\right\rVert\leq\bar{\sigma}$ for all $t\geq 0$.
@@
These assumptions are standard in stochastic optimization literature bottou2018optimization ; ghadimi2013stochastic ; gower2020variance , and they hold naturally or can be easily enforced for typical neural network architectures and loss functions used in machine learning practice. Notably, the variance of the stochastic gradient estimator (Assumption 5.2) is more directly controllable from a practitioner’s perspective by selecting a sufficiently large batch size or adding a momentum term that uses past gradients to inform the direction of update (see, e.g., bottou2018optimization ; gower2020variance ; defazio2019ineffectiveness ). We also remark that the matrix $\mathbf{B}_{t}$ in Assumption 5.4 may be interpreted as a preconditioner. Depending on the nature of its conditional correlation with $\mathbf{g}_{t}$ given $\mathbf{w}_{t}$, a large batch size can enhance the practicality of the assumption. In practice, either this conditional correlation is nonexistent (see the comment on (bottou2018optimization, , Assumption 4.3(b))), or the contribution of $\mathbf{g}_{t}$ is inversely scaled by the (large) batch size as in the proof of Lemma 5.5 below.
In Lemma 5.5, we prove a descent lemma for EGN under the given conditions.
@@
Let $\{\mathbf{w}_{t}\}$ be the sequence of iterates generated by (3.4) with $\mathbf{d}_{t}\equiv\mathbf{d}_{t}^{\mathrm{LM}}$ and let Assumptions 5.15.3 hold. Suppose there exists $\alpha^{\text{max}}>0$ such that $\alpha_{t}\leq\alpha^{\text{max}}$ for all $t$ in Algorithm 4. Then,
@@
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\frac{\alpha_{t}}{2}\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{3\sigma_{g}^{2}\alpha_{t}^{2}}{10b\lambda_{t}\alpha^{\text{max}}}.$ (5.6)
Proof. From Assumption 5.1, we have
$={\mathcal{L}}_{N}(\mathbf{w}_{t})-\alpha_{t}\left\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}\mathbf{g}_{t}\right\rangle+\frac{\alpha_{t}^{2}L_{1}}{2}\left\lVert\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}\mathbf{g}_{t}\right\rVert^{2}.$ (5.7)
Next, we set $\mathbf{B}_{t}=\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}$ in Assumption 5.4. From Assumption 5.3, we can show that $\mathbf{B}_{t}\succeq\mu\mathbf{I}$ with $\mu\coloneqq b/(b\lambda_{t}+\bar{\kappa}\bar{\sigma}^{2})$. Using this result and Assumption 5.3 in (5.7), we obtain
${\mathcal{L}}_{N}(\mathbf{w}_{t+1})$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\alpha_{t}\left\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{B}_{t}\mathbf{g}_{t}\right\rangle+\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}\|\mathbf{g}_{t}\|^{2}.$ (5.8)
Notice that by Assumption 5.2, we have ${\mathbb{E}}[\mathbf{g}_{t}|\mathbf{w}_{t}]=\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})$. Hence, the inequality in Assumption 5.4 can be written as ${\mathbb{E}}\left[\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{B}_{t}\mathbf{g}_{t}\rangle|\mathbf{w}_{t}\right]\geq\mu K\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}$. Taking conditional expectation on both sides of (5.8) with respect to $\xi$ and using Assumption 5.4, we get
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\frac{3\alpha_{t}}{5\lambda_{t}}\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}{\mathbb{E}}[\|\mathbf{g}_{t}\|^{2}|\mathbf{w}_{t}].$ (5.9)
Using Assumption 5.2, we have
$={\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]$
$\qquad+2{\mathbb{E}}[\langle\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\rangle|\mathbf{w}_{t}]+{\mathbb{E}}[\|\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]$
$\leq\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{\sigma_{g}^{2}}{b}.$ (5.10)
Now, using (5.10) in (5.9), we obtain
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\left(\frac{3\alpha_{t}}{5\lambda_{t}}-\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}\right)\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{b\sigma_{g}^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}.$ (5.11)
By choosing $\alpha_{t}$ to be sufficiently small, we can set $\alpha^{\text{max}}=1/c_{1}$ in Algorithm 4, where $c_{1}=(5b^{2}\lambda_{t}L_{1})/(3(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}$. With this selection, inequality (5.11) holds, which completes the proof.
$\Box$
@@
Assumption 5.4 is necessary to prove the result in Lemma 5.5 due to the correlation between $\mathbf{B}_{t}$ and $\mathbf{g}_{t}$. If these quantities were not correlated, then one could exploit Assumptions 5.25.3: by taking the expected value and using (5.4), one could then bound $\mathbf{B}_{t}\succeq\mu\mathbf{I}$ to obtain the desired bound with $K=1$. Clearly, that would yield different constants in our next results, but their nature would remain unaltered. Note that this suggests alternative versions of EGN, in which the correlation is removed by construction at the expense of an increased computational burden. A thorough investigation of such schemes is beyond the scope of this paper.
Unlike classical stochastic quasi-Newton and SGD ($\mathbf{H}_{t}=\mathbf{I}_{d}$) methods, the loss decrease condition of EGN has an explicit nonlinear dependence on the batch size $b$. In the regime $b\gg 1$, the influence of the variance parameter $\sigma_{g}$ diminishes.
We prove next the following result about the convergence of EGN.